Find the first and second derivatives of the following and simplify whenever possible:



x = a cosh t; y=b sinh t

1
Expert's answer
2021-11-30T06:13:09-0500

"y'_x = \dfrac{y'_t} {x'_t} = \dfrac{b \cosh t}{a\sinh t} = \dfrac{b} {a\tanh t}, \\
\\
y''_{xx} = \dfrac{(y' _x) '_t} {x'_t} = \dfrac{\frac{b}{a}\cdot \dfrac{\sinh^2t-\cosh^2t}{\sinh^2t} } {a\sinh t} = - \dfrac{b} {a^2\sinh^3t}."